OpenStudy (anonymous):

Boris was 26 years old in 1970, when his daughter, Olga, was born. In what year was Boris exactly 3 times as old as Olga?

OpenStudy (lalaly):

boris age=B olgas age = C (B+x)=3(C+x) B=26 C=0 26+x=0+3x 26=2x x=13 so 1970+13=1983